An unmanned vehicle follows a curved path, beginning at O and ending at A, of length 18 m. At O the speed of the vehicle is 12 m/s and it has slowed to 6 m/s at A. The tangential deceleration of the vehicle is proportional to the distance from O. a) Write an equation for tangential acceleration. b) Find the constant of proportionality in your equation


1
Expert's answer
2022-05-12T15:09:13-0400

"a=\frac{v_1^2-v_2^2}r=-6~\frac m{s^2}."